# A Small Collusion is All You Need

### v1

@@ Abstract.
Transaction Fee Mechanisms (TFMs) study auction design in the Blockchain context, and emphasize robustness against miner and user collusion, moreso than traditional auction theory. (Chung and Shi, 2023) introduce the notion of a mechanism being $c$-Side-Contract-Proof ($c$-SCP), i.e., robust to a collusion of the miner and $c$ users. Later work (Chung et al., 2024; Gafni and Yaish, 2025) shows a gap between the $1$-SCP and $2$-SCP classes. We show that the class of $2$-SCP mechanisms equals that of any $c$-SCP with $c\geq 2$, under a relatively minor assumption of consistent tie-breaking. In essence, this implies that any mechanism vulnerable to collusion, is also vulnerable to a small collusion.
@@ 1. Introduction
Transaction Fee Mechanisms (TFMs) is the name given to the class of mechanisms responsible for the ordinary operation of a blockchain system, choosing which transactions are confirmed at each issued block. In Bitcoin, this is done by having each transaction commit to a fee that it pays if confirmed. This is in essence a pay-as-bid auction, and a line of work since (Lavi et al., 2019; Huberman et al., 2021) has considered alternative auction designs. Importantly, this has not remained a theoretical exercise, and major blockchain protocols have adopted different mechanisms, most notably Ethereum, which adopted the EIP-1559 mechanism (in essence, a first-price with a reserve that is fully burned) (Roughgarden, 2020a).
@@
TFMs have several unique characteristics that distinguish them from traditional auctions (Tim Roughgarden, 2024): (i) They allow for payments to be “burned”, i.e., the user pays but the miner (which takes the role of the auctioneer) does not receive the full payment, (ii) Their research is more focused on robustness to different forms of attacks, rather than just incentive-compatibility w.r.t. the bidders, and welfare/revenue maximization within this framework, which is the standard classic approach to auctions (Vickrey, 1961; Myerson, 1981). In particular, the issue of collusion-resistance has taken an important role in TFM research (Roughgarden, 2020b; Chung and Shi, 2023). Even in traditional auctions, practitioners have long emphasized the importance of collusion-resistance. The following is an emblematic quote:
@@
“The most important features of an auction are its robustness against collusion and its attractiveness to potential bidders. Failure to attend to these issues can lead to disaster.” (Klemperer, 2002)
@@
However, this takes maybe even greater precedence in Blockchain systems, because of their online and ephemeral nature. In an anonymous, unregulated environment, bidders may worry less about facing prosecution for cartel behavior, as they have faced many times in traditional auctions (Posner, 1970). Thus, TFM research strives to make collusion-resistance a built-in part of its design.
@@
We now arrive at the topic at the center of this paper. (Gafni and Yaish, 2022) raise the question, supported by influential work in economic theory (Stigler, 1964), of whether collusion is a real risk in TFMs. The argument goes, that since in some systems block times are quick (e.g., $12$ seconds in Ethereum), it is hard to not only establish a collusion, but also coordinate, through possibly several rounds of communication, how each agent should partake in it. Both in traditional auctions and in TFMs, there is some support in theory and empirics that the smaller the collusion, the easier it is to establish it, coordinate it, and avoid its detection. Thus, it is important to know whether mechanisms are robust to small-size collusion.
@@
Our work does exactly that: We consider the popular collusion notion of $c$-SCP, introduced by (Chung and Shi, 2023). The notion considers whether a coalition of the miner and up to $c$ bidders is able to increase its utility by changing bids, as well as using manipulations by the miner (dropping bids, or adding shill bids). The $c$ parameter naturally establishes a hierarchy: Clearly, the class of $1$-SCP mechanisms (i.e., these robust to a collusion of the miner and at most a single user) is the widest class, while the class of SCP (used as shorthand for $n$-SCP, i.e., a collusion of the miner and any number of bidders) is the narrowest. Moreover, an example of (Chung et al., 2024) shows a gap between the class of $1$-SCP and $2$-SCP mechanisms. It is thus natural to ask,
@@ 1.1.1. The Core TFM Framework and Extensions
Considering blockchain transaction fee mechanisms as an auction design problem was studied by (Lavi et al., 2019; Huberman et al., 2021). The early emphasis was on miner incentive-compatibility (MIC), i.e., robustness to manipulations by the miner such as dropping bids, and introducing fake bids. (Roughgarden, 2020b) was the first to introduce collusion-resilience as the centerpiece of TFM design, through the notion of OCA-proof. OCA-proof is a notion that requires the mechanism to be welfare-maximizing for the grand coalition. (Chung and Shi, 2023) introduce the $c$-SCP notion, which captures a more game-theoretic group deviation notion, which we study in this paper. The focus of the axiomatic literature that followed was on showing whether a “dream” TFM, that combines collusion-resilience, as well as the user and miner incentive-compatibility notions, exists. The overall conclusion is that the answer is no (To mention some results, no user incentive-compatible (UIC) + $1$-SCP mechanism (Chung and Shi, 2023) exists, no UIC+MIC+OCA-proof deterministic mechanism, scalable randomized one, or randomized where the collusion knows the random coins (Chung et al., 2024; Gafni and Yaish, 2024a) exists), although some open questions remain. In particular, a non-scalable randomized mechanism where the collusion is oblivious to the random coins may still exist, though its welfare is bounded away from optimal (Gafni and Yaish, 2024a). However, one might be interested in a subset of the restrictions. For example, if we think that regular users are ephemeral and have bounded rationality, we may omit the user incentive-compatibility requirement. One reason to think so is the common use of fee estimators and calculators by users (Zhang et al., 2024)). With only a subset of restrictions, there are characterizations the classes of mechanisms that implement them (Gafni and Yaish, 2024a). Importantly, we are not aware of a past work that examined the notion of SCP in isolation, and drew conclusions and characterizations from it alone, and this is a novelty of our work.
Beyond the core framework, the study of TFMs was extended in many directions: (1) Considering time preferences (Gafni and Yaish, 2024b; Nisan, 2023; Babaioff and Nisan, 2024), (2) Considering the problem of Maximal Extractable Value (MEV), i.e., incentive problems that may arise in certain blockchain use-cases where the order of the transactions in the block matters (Bahrani et al., 2024), (3) Considering more advanced blockchain architectures, such as DAGs and PBS (Bahrani et al., 2025; Garimidi et al., 2025). We emphasize that our work is in the core TFM setting, and does not consider these variations.
@@ 1.2. An Overview of Our Proof
Our proof is constructive and presents a reduction from a given general collusion to a $2$-collusion. First, we make sure the collusion is in what we call “Canonical Ordering”. This means that bidders in the collusion do not change their relative order before and after the collusion. This is possible since a collusion can always have even better utility by allowing higher-value agents to be confirmed. Keeping the canonical ordering throughout our reduction is useful in making the problem one-dimensional (’how many colluding bidders are confirmed?’) instead of multi-dimensional (’is each of the colluding bidders confirmed?’). The main idea is to decompose the collusion into steps, where at each step only one bidder changes their bid. The goal is to isolate one step which is in itself a beneficial collusion. However, this has to be done carefully. Two important insights allow the decomposition:
@@ 2. Model and Preliminaries
We follow the standard model used by the literature for Transaction Fee Mechanisms (TFMs) (Lavi et al., 2019; Roughgarden, 2024; Chung and Shi, 2023; Shi et al., 2023). The model assumes a multi-unit, unit-demand setting, i.e., bidders have a fixed-size transaction, and do not care about specific placement in the block. We interchangeably use the terms auction, mechanism, or transaction fee mechanism (TFM)to refer to the object of our discussion, depending on what is most appropriate in the context.
@@ 2.1. Transaction Fee Mechanisms
Lastly, we introduce the notion of zero-utility bidders and make the following assumption:
@@
Zero-utility bidders. We call a bidder a zero-utility bidder in a setting $A$ if it is either unconfirmed, or is confirmed and pays their bid.
Consistent Tie-Breaking towards zero-utility bidders w.r.t. unconfirmed bidders.
@@
This definition is natural in the sense that it holds for reasonable auction formats, such as posted-price, or second-price and first-price auctions, with or without reserve. It also holds for weirder specially-constructed formats, such as the discount auction of (Chung et al., 2024).
@@ 3. The class of $2$-SCP mechanisms equals the class of SCP mechanisms
As observed by (Gafni and Yaish, 2025), the Discount auction of (Chung et al., 2024) shows a nice separation between $1$-SCP and $2$-SCP. We show that there are no other gaps between $c$-SCP and $(c+1)$-SCP mechanisms. We start by noting that we can restrict ourselves to the passive miner model of collusion.
@@
Consider if for some two bids $b_{i}>b_{j}$, the latter is confirmed while the former is not, and consider the $2$-SCP where bidder $i$ bids $b_{j}$ and bidder $j$ bids $b_{i}$. The miner utility is unchanged, the sum of utilities for agents $i,j$ increases by $b_{i}-b_{j}$, so it is a $2$-SC. ∎
@@
Figure 1. A demonstration of Lemma $3.$ Given a collusion where bidders change their order after the collusion, we can find a collusion where the order is maintained.: A demonstration of Lemma $3.$ Given a collusion where bidders change their order after the collusion, we can find a collusion where the order is maintained.
@@
Consider a bidder $i$ bids $b_{i}$ in $A$ and $b^{\prime}_{i}$ in $B$, and this is the only change in bids between the settings. Assume bidder $i$ is a zero-utility bidder in both settings, with the same confirmation status in both. Then, the number of non zero-utility bidders in $B$ is the same as in $A$.
@@
Assume towards contradiction there is a bidder $j$ that is a zero-utility bidder in $A$ and a non-zero-utility bidder in $B$. Consider a collusion between the miner and bidders $i,j$. Then, by the $2$-SCP conditions from $A$ to $B$ and vice versa,
$\begin{split}&u_{miner}(A)=u_{miner}(A)+u_{i}(A;A)+u_{j}(A;A)\\ &\geq u_{miner}(B)+u_{i}(B;A)+u_{j}(B;A),\\ &u_{miner}(B)=u_{miner}(B)+u_{i}(B;B)+u_{j}(B;B)\\ &\geq u_{miner}(A)+u_{i}(A;B)+u_{j}(B;A),\end{split}$
@@
(6) $\begin{split}&u_{miner}(A)\geq u_{miner}(A)+u_{i}(A;B)+u_{i}(B;A)+u_{j}(B;A)+u_{j}(A;B)\end{split}$
For bidder $j$, since it does not change its bid between $A$ and $B$, and it is a zero-utility bidder in $A$ and non-zero-utility bidder in $B$, it holds that:
$u_{j}(A;B)+u_{j}(B;A)=0+b_{j}-p_{i}^{B}>0.$
For bidder $i$, if it is confirmed in both $A$ and $B$, then:
$u_{i}(A;B)+u_{i}(B;A)=b^{\prime}_{i}-p_{i}^{A}+b_{i}-p_{i}^{B}=0,$
and if it is unconfirmed in both $A$ and $B$, we also have $u_{i}(A;B)+u_{i}(B;A)=0$.
Thus, in both cases we have
$u_{miner}(A)+u_{i}(A;B)+u_{i}(B;A)+u_{j}(B;A)+u_{j}(A;B)>u_{miner}(A),$
which is in contradiction to Eq. 6.
@@
Figure 2. The “Salsa Dance” Lemma. We separate the bidders into $U_{I},D,U_{O}$, and let each bidder change their bid one-by-one following Definition 3.6. Lemma 3.8 restricts the possible confirmations of each step in the decomposition, and the figure shows one such valid sequence of confirmations.: The “Salsa Dance” Lemma. We separate the bidders into $U_{I},D,U_{O}$, and let each bidder change their bid one-by-one following Definition 3.6. Lemma 3.8 restricts the possible confirmations of each step in the decomposition, and the figure shows one such valid sequence of confirmations.
@@
The proof of ()(*) is a matter of case-analysis and we show it in the appendix. We conclude that Eq. 12 shows $0\geq\delta$, but since $\delta>0$, this is a contradiction, and there must be some two subsequent steps constitute a beneficial collusion. ∎

### v2

@@ Abstract.
Transaction Fee Mechanisms (TFMs) study auction design in the Blockchain context, and emphasize robustness against miner and user collusion, moreso than traditional auction theory. (chung2023foundations) introduce the notion of a mechanism being $c$-Side-Contract-Proof ($c$-SCP), i.e., robust to a collusion of the miner and $c$ users. Later work (chung2024collusion; welfareIncreasingCollusion) shows a gap between the $1$-SCP and $2$-SCP classes. We show that the class of $2$-SCP mechanisms equals that of any $c$-SCP with $c\geq 2$, under a relatively minor assumption of consistent tie-breaking. In essence, this implies that any mechanism vulnerable to collusion, is also vulnerable to a small collusion.
@@ 1. Introduction
Transaction Fee Mechanisms (TFMs) is the name given to the class of mechanisms responsible for the ordinary operation of a blockchain system, choosing which transactions are confirmed at each issued block. In Bitcoin, this is done by having each transaction commit to a fee that it pays if confirmed. This is in essence a pay-as-bid auction, and a line of work since (lavi2019redesigning; huberman2021monopoly) has considered alternative auction designs. Importantly, this has not remained a theoretical exercise, and major blockchain protocols have adopted different mechanisms, most notably Ethereum, which adopted the EIP-1559 mechanism (in essence, a first-price with a reserve that is fully burned) (roughgarden2020eip1559).
@@
TFMs have several unique characteristics that distinguish them from traditional auctions (roughgarden8reasons): (i) They allow for payments to be “burned”, i.e., the user pays but the miner (which takes the role of the auctioneer) does not receive the full payment, (ii) Their research is more focused on robustness to different forms of attacks, rather than just incentive-compatibility w.r.t. the bidders, and welfare/revenue maximization within this framework, which is the standard classic approach to auctions (vickrey1961; myerson1981optimal). In particular, the issue of collusion-resistance has taken an important role in TFM research (roughgarden2020transaction; chung2023foundations). Even in traditional auctions, practitioners have long emphasized the importance of collusion-resistance. The following is an emblematic quote:
@@
“The most important features of an auction are its robustness against collusion and its attractiveness to potential bidders. Failure to attend to these issues can lead to disaster.” (klempererWhatMatters)
@@
However, this takes maybe even greater precedence in Blockchain systems, because of their online and ephemeral nature. In an anonymous, unregulated environment, bidders may worry less about facing prosecution for cartel behavior, as they have faced many times in traditional auctions (posner70). Thus, TFM research strives to make collusion-resistance a built-in part of its design.
@@
We now arrive at the topic at the center of this paper. (gafni2022greedy) raise the question, supported by influential work in economic theory (Stigler1964), of whether collusion is a real risk in TFMs. The argument goes, that since in some systems block times are quick (e.g., $12$ seconds in Ethereum), it is hard to not only establish a collusion, but also coordinate, through possibly several rounds of communication, how each agent should partake in it. Both in traditional auctions and in TFMs, there is some support in theory and empirics that the smaller the collusion, the easier it is to establish it, coordinate it, and avoid its detection. Thus, it is important to know whether mechanisms are robust to small-size collusion.
@@
Our work does exactly that: We consider the popular collusion notion of $c$-SCP, introduced by (chung2023foundations). The notion considers whether a coalition of the miner and up to $c$ bidders is able to increase its utility by changing bids, as well as using manipulations by the miner (dropping bids, or adding shill bids). The $c$ parameter naturally establishes a hierarchy: Clearly, the class of $1$-SCP mechanisms (i.e., these robust to a collusion of the miner and at most a single user) is the widest class, while the class of SCP (used as shorthand for $n$-SCP, i.e., a collusion of the miner and any number of bidders) is the narrowest. Moreover, an example of (chung2024collusion) shows a gap between the class of $1$-SCP and $2$-SCP mechanisms. It is thus natural to ask,
@@ 1.1.1. The Core TFM Framework and Extensions
Considering blockchain transaction fee mechanisms as an auction design problem was studied by (lavi2019redesigning; huberman2021monopoly). The early emphasis was on miner incentive-compatibility (MIC), i.e., robustness to manipulations by the miner such as dropping bids, and introducing fake bids. (roughgarden2020transaction) was the first to introduce collusion-resilience as the centerpiece of TFM design, through the notion of OCA-proof. OCA-proof is a notion that requires the mechanism to be welfare-maximizing for the grand coalition. (chung2023foundations) introduce the $c$-SCP notion, which captures a more game-theoretic group deviation notion, which we study in this paper. The focus of the axiomatic literature that followed was on showing whether a “dream” TFM, that combines collusion-resilience, as well as the user and miner incentive-compatibility notions, exists. The overall conclusion is that the answer is no (To mention some results, no user incentive-compatible (UIC) + $1$-SCP mechanism (chung2023foundations) exists, no UIC+MIC+OCA-proof deterministic mechanism, scalable randomized one, or randomized where the collusion knows the random coins (chung2024collusion; gafni2024barriers) exists), although some open questions remain. In particular, a non-scalable randomized mechanism where the collusion is oblivious to the random coins may still exist, though its welfare is bounded away from optimal (gafni2024barriers). However, one might be interested in a subset of the restrictions. For example, if we think that regular users are ephemeral and have bounded rationality, we may omit the user incentive-compatibility requirement. One reason to think so is the common use of fee estimators and calculators by users (zhang2024transactionfeeestimationbitcoin)). With only a subset of restrictions, there are characterizations of the classes of mechanisms that implement them (gafni2024barriers). Importantly, we are not aware of a past work that examined the notion of SCP in isolation, and drew conclusions and characterizations from it alone, and this is a novelty of our work.
Beyond the core framework, the study of TFMs was extended in many directions: (1) Considering time preferences (gafni2024scheduling; nisan2023serial; babaioff2024optimality), (2) Considering the problem of Maximal Extractable Value (MEV), i.e., incentive problems that may arise in certain blockchain use-cases where the order of the transactions in the block matters (bahrani2024transaction), (3) Considering more advanced blockchain architectures, such as DAGs and PBS (garimidiPBS; garimidi2025dag). We emphasize that our work is in the core TFM setting, and does not consider these variations.
@@ 1.2. An Overview of Our Proof
Our proof is constructive and presents a reduction from a given general collusion to a $2$-collusion. First, we make sure the collusion is in what we call “Canonical Ordering”. This means that bidders in the collusion do not change their relative order, in terms of value, before and after the collusion. This is possible since a collusion can always have even better utility by allowing higher-value agents to be confirmed. Keeping the canonical ordering throughout our reduction is useful in making the problem one-dimensional (’how many colluding bidders are confirmed?’) instead of multi-dimensional (’is each of the colluding bidders confirmed?’). The main idea is to decompose the collusion into steps, where at each step only one bidder changes their bid. The goal is to isolate one step which is in itself a beneficial collusion. However, this has to be done carefully. Two important insights allow the decomposition:
@@ 2. Model and Preliminaries
We follow the standard model used by the literature for Transaction Fee Mechanisms (TFMs) (lavi2019redesigning; roughgarden2024transaction; chung2023foundations; shi2023what). The model assumes a multi-unit, unit-demand setting, i.e., bidders have a fixed-size transaction, and do not care about specific placement in the block. We interchangeably use the terms auction, mechanism, or transaction fee mechanism (TFM)to refer to the object of our discussion, depending on what is most appropriate in the context.
@@ 2.1. Transaction Fee Mechanisms
We say a bidder is a zero-utility bidder if $u_{i}(b_{i},\mathbf{b_{-i}};b_{i})=0$: I.e., it is a bidder that is either unconfirmed, or pays their bid.
We make the following tie-breaking assumption:
@@
Consistent Tie-Breaking towards zero-utility bidders, w.r.t. unconfirmed bidders.
@@
This definition is natural in the sense that it holds for reasonable auction formats.
Example 2.1.
Consider a single-item posted-price auction with price $p$ and lexicographic tie-breaking, i.e., for bidders $b_{1},\ldots,b_{n}$, the lowest index bidder with $b_{i}\geq p$ wins the item and pays $p$. Then this tie-breaking is consistent towards zero-utility bidders, w.r.t. unconfirmed bidders. Let us unpack the definition directly: Consider some unconfirmed bidder $j$. It either has a higher index than the confirmed bidder, or $b_{j}<p$. Consider it changes its bid and remains unconfirmed (then, in the latter case, we have $b^{\prime}_{j}<p$). Now consider a zero-utility bidder $i$: If $i$ is unconfirmed, then $i$ remains unconfirmed. If $i$ is confirmed, then it must be that $b_{i}=p$, and it remains confirmed in both settings.
Alternatively, consider a second-price auction with lexicographic tie-breaking. I.e., the highest bidder wins, and if there are several equal highest bids, the lowest indexed one wins. Similarly, this is consistent towards zero-utility bidders, w.r.t. unconfirmed bidders.
We note that the property holds for specially-constructed formats, such as the discount auction of (chung2024collusion).
@@ 3. The class of $2$-SCP mechanisms equals the class of SCP mechanisms
As observed by (welfareIncreasingCollusion), the Discount auction of (chung2024collusion) shows a nice separation between $1$-SCP and $2$-SCP. We show that there are no other gaps between $c$-SCP and $(c+1)$-SCP mechanisms. We start by noting that we can restrict ourselves to the passive miner model of collusion.
@@
Consider if for some two bids $b_{i}>b_{j}$, the latter is confirmed while the former is not, and consider the $2$-SC where bidder $i$ bids $b_{j}$ and bidder $j$ bids $b_{i}$. The miner utility is unchanged, the sum of utilities for agents $i,j$ increases by $b_{i}-b_{j}$, so it is a $2$-SC. ∎
@@
Figure 1. A demonstration of Lemma $3.$ Given a collusion where bidders change their order after the collusion, we can find a collusion where the order is maintained.
[…]
@@
Consider a bidder $i$ bidding $b_{i}$ in $A$ and $b^{\prime}_{i}$ in $B$, and this is the only change in bids between the settings. Assume bidder $i$ is unconfirmed in both settings. Then, the number of non zero-utility bidders in $B$ is the same as in $A$.
@@
Assume towards contradiction there is a bidder $j$ that is a zero-utility bidder in $A$ and a non-zero-utility bidder in $B$.
First, consider a collusion between the miner and bidder $i$. Then, by the $1$-SCP condition from $A$ to $B$ and vice versa,
$\begin{split}&u_{miner}(A)=u_{miner}(A)+u_{i}(A;A)\\ &\geq u_{miner}(B)+u_{i}(B;A)=u_{miner}(B),\\ &u_{miner}(B)=u_{miner}(B)+u_{i}(B;B)\\ &\geq u_{miner}(A)+u_{i}(A;B)=u_{miner}(A),\end{split}$
which yields $u_{miner}(A)=u_{miner}(B)$.
Next, consider a collusion between the miner and bidders $i,j$. Then, by the $2$-SCP conditions from $A$ to $B$ and vice versa,
$\begin{split}&u_{miner}(A)=u_{miner}(A)+u_{i}(A;A)+u_{j}(A;A)\\ &\geq u_{miner}(B)+u_{i}(B;A)+u_{j}(B;A),\\ &u_{miner}(B)=u_{miner}(B)+u_{i}(B;B)+u_{j}(B;B)\\ &\geq u_{miner}(A)+u_{i}(A;B)+u_{j}(A;B),\end{split}$
@@
(6) $\begin{split}&u_{miner}(A)\geq u_{miner}(A)+u_{j}(A;B)+u_{j}(B;A),\end{split}$
i.e.,
$u_{j}(A;B)+u_{j}(B;A)\leq 0.$
However, notice that for bidder $j$, since it does not change its bid between $A$ and $B$, and it is a zero-utility bidder in $A$ and non-zero-utility bidder in $B$, it holds that:
$u_{j}(A;B)+u_{j}(B;A)=0+b_{j}-p_{i}^{B}>0,$
in contradiction. ∎
@@
Figure 2. The “Salsa Dance” Lemma. We separate the bidders into $U_{I},D,U_{O}$, and let each bidder change their bid one-by-one following Definition 3.6. Lemma 3.9 restricts the possible confirmations of each step in the decomposition, and the figure shows one such valid sequence of confirmations.
[…]
Before we prove the main structural lemma for the “Salsa Dance” decomposition, we demonstrate that the specific choice of decomposition matters for its success.
Example 3.8.
Consider an auction with $k=2$ items with the allocation rule that the two highest bidders $v_{1},v_{2}$ are both confirmed if and only if $v_{1}+v_{2}\geq 2$. Regarding payments, $v_{2}$ pays their bid, and $v_{1}$ pays the bids’ average $\frac{v_{1}+v_{2}}{2}$. All payments are burned. Consider bidders with true values $b_{1}=1,b_{2}=1$ that bid $b^{\prime}_{1}=\frac{5}{4},b^{\prime}_{2}=\frac{3}{4}$. This is a $2$-SC that brings the joint utility of the two bidders and the miner from 0 $\frac{1}{4}$. Under the Salsa Decomposition, the two steps are $(1,1)\rightarrow(5/4,1)\rightarrow(5/4,3/4)$, and the second step is a $2$-SC (the joint utility changes from $\frac{1}{8}$ to $\frac{1}{4}$ (recall that for each decomposed step, we consider the beginning of the step to be the true values and the end of the step to be the collusion bids). The other possible decomposition is $(1,1)\rightarrow(1,3/4)\rightarrow(5/4,3/4)$, where no step is a $2$-SC.
@@
The proof of ()(*) is a matter of case-analysis and we show it in the appendix. We conclude that Eq. 12 shows $0\geq\delta$, but since $\delta>0$, this is a contradiction, and there must be some two subsequent steps constituting a beneficial collusion. ∎